OpenStudy (anonymous):

A shirt on sale for $ 15.00 and has been marked down 35%. how much was the shirt being sold for before the sale?

OpenStudy (mathteacher1729):

If the old shirt cost 100 dollars, and is then marked down 35%, that means it would cost $65 after markdown. (notice that in decimals1.00 - 0.35 = 0.65) The problem here is that the shirt did not originally cost $100 or that would be too easy. :-p Since it costs $15 AFTER being marked down... we must solve for the old price (call it P) (old price) times ( 0.65) = 15 or 0.65P = 15 Hope this helps. :)

OpenStudy (anonymous):

thanks ^_^ it did help!

OpenStudy (anonymous):

so p= 15.00/0.65=23.0769.....

OpenStudy (mathteacher1729):

Two things about your answer: 1) Test it to make sure it's true! (take P and multiply it by 0.65, which is the same as taking off 35%). 2) Dollars and Cents only go up to the hundredths place, so round your answer accordingly. Nobody pays a bill for $23.0769! ;-p Good work!

OpenStudy (anonymous):

LOL ok 23.08!!

OpenStudy (sasogeek):

after working this, i had 42.86 if 15.00 is 35% of the original price, then the original price is unknown so say, Y. 35% x Y=15 => Y=1500/35 Y=42.86 approximately

OpenStudy (mathteacher1729):

Saosogeek -- sorry but that is incorrect. To take off 35% means: (original price) MINUS (35% of original price) = (new price) If we let P = original price and $15 be the new price: P - 0.35P = 15 But the left side is just 0.65P so that is why you write 0.65P = 15 The answer MekMek has is correct.

OpenStudy (sasogeek):

i think i get it now :) thanks for the further explanation :)